{L \atop l}

L l m7.ppt” ><

INP UT T Y P E = ”image”NAME =

”LaunchP ASCOinterfaceapplication”SRC = ”../gif/telescope.gif”ALIGN =

”MIDDLE” >< INP UT T Y P E = ”submit”V ALUE = ”LaunchP ASCOapplication” ><

/F ORM >< /T D >< /T R >< /T ABLE > button.epsicon.Letthebobswingforabout17periods.Calculatethemeanandstandarddeviationbysimplyclickingonthetablestatisticsicon(i.e. Σ ).F orasinglemeasurementoftheperiodthestandarddeviationisareasonablemeasureoftheuncertainty.W ith17measurementstheuncertaintyofthemeanisσ/\sqrt{N − 1}

N = 17 ?? L L ?? ∘

Plot the measured period as a function of angular amplitude including a few error bars. θ

T = T_{0} (1 + \frac{1}{2^{2}} sin^{2} \frac{θ}{2} + \frac{1(3^{2})}{2^{2}(4^{2})} sin^{4} \frac{θ}{2} + ⋯)

T_{0} = 2π\sqrt{(L/g)} θ

θT/T_{0}

Period vs. Length:

T T_{0} T_{0} L L = 0 T_{0} L T_{0}^{2}

T_{0} L Measurement of T = 2π\sqrt{(L/g)} T L g Calculate the uncertainty in your determination of g. Δg = g\sqrt{(ΔL/L)^{2} + (2ΔT /T)^{2}}

M = 62 L_{1}+

L_{2} L_{2}

T = T_{0} (1 + \frac{ρ}{2ρ} )

T_{0} ρ T

m

T = T_{0} (1 − \frac{m}{12M} )

M g

L_{1}+ L_{2}